need answer for this question

Ferrier Chemical Company makes three products, B7, K6, and X9, which are joint products from the same materials. In a standard batch of 150,000 pounds of raw materials, the company generates 35,000 pounds of B7, 75,000 pounds of K6, and 40,000 pounds of X9. A standard batch costs $600,000 to produce. The sales prices per pound are $6, $10, and $16 for B7, K6, and X9, respectively.

Required

Allocate the joint product cost among the three final products using weight as the allocation base.

Allocate the joint product cost among the three final products using market value as the allocation base.
Answer #1

(1): Using weight as the allocation base: Here the formula is cost allocated to a joint product = weight of the product/total weight*total joint costs

Weight of the product Total weight Allocation rate = weight of the product/total weight Joint product cost = allocation rate*cost of $600,000
B7 35,000.00 150,000.00 0.2333 140,000.00
K6 75,000.00 0.5000 300,000.00
X9 40,000.00 0.2667 160,000.00
Total 600,000.00

(2): Market value as allocation base:

Here the allocation basis = sale value of the product/sale value of total production

Sale value of B7 = $6*35,000 = 210,000

Sale value of K6 = $10*75,000 = 750,000

Sale value of X9 = $16*40,000 = 640,000.

Thus total sale value = 210,000+750,000+640,000 = $1,600,000

Now joint costs for B7 = 210,000/1,600,000*$600,000 = 78,750

Sale value Total sale Allocation rate = sale value/total sales Joint product cost = allocation rate*cost of $600,000
B7 210,000.00 1,600,000.00 0.1313 78,750.00
K6 750,000.00 0.4688 281,250.00
X9 640,000.00 0.4000 240,000.00
Total 600,000.00